### v1

# One-to-one correspondence reconstruction at the electron-positron Higgs factory

Abstract
We propose one-to-one correspondence reconstruction for electron-positron Higgs factories. For each visible particle, one-to-one correspondence aims to associate relevant detector hits with only one reconstructed particle and accurately identify its species. To achieve this goal, we develop a novel detector concept featuring 5-dimensional calorimetry that provides spatial, energy, and time measurements for each hit, and a reconstruction framework that combines state-of-the-art particle flow and artificial intelligence algorithms. In the benchmark process of Higgs to di-jets, over 90% of visible energy can be successfully mapped into well-reconstructed particles that not only maintain a one-to-one correspondence relationship but also associate with the correct combination of cluster and track, improving the invariant mass resolution of hadronically decayed Higgs bosons by 25%. Performing simultaneous identification on these well-reconstructed particles, we observe efficiencies of 97% to nearly 100% for charged particles ($e^{\pm}$, $\mu^{\pm}$, $\pi^{\pm}$, $K^{\pm}$, $p/\bar{p}$) and photons ($\gamma$), and 75% to 80% for neutral hadrons ($K_{L}^{0}$, $n$, $\bar{n}$). For physics measurements of Higgs to invisible and exotic decays, golden channels to probe new physics, one-to-one correspondence could enhance discovery power by 10% to up to a factor of two. This study demonstrates the necessity and feasibility of one-to-one correspondence reconstruction at electron-positron Higgs factories.
In collider experiments, high-energy particles impact each other and generate multiple final state particles, bringing the information of underlying physics laws. These particles interact with the detector, excite detector sensors, and generate electronic signals that are recorded into data, typically in the form of detector hits. Correspondingly, data processing in collider experiments consists of two steps: interpreting detector hits into physics objects, especially final state particles, and performing physics measurements using these physics objects. The first step is called reconstruction, which can be regarded as the inverse process of detector hits generation.
The reconstruction establishes a mapping between visible particles and reconstructed ones, where the ultimate goal is a one-to-one (1-1) correspondence. The visible particles include both primary particles from the interaction point (IP) and secondary particles generated through interactions with detector materials upstream of the calorimeter, as far as they create detector hits. 1-1 correspondence reconstruction aims to not only successfully reconstruct each individual particle, but also correctly identify its species (a.k.a. the particle identification, PID). 1-1 correspondence is highly beneficial for collider experiments because it provides a universal foundation for reconstructing various physics objects and can significantly enhance the reconstruction performance for jets and missing energy/momentum.
Depending on the multiplicity of visible particles and the detector configuration, achieving 1-1 correspondence could be truly challenging. For proton-proton and heavy ion collisions at the LHC [1] or future hadron colliders, each collision can generate $\mathcal{O}(10^{2}\text{--}10^{4})$ visible particles, making 1-1 correspondence extremely difficult, especially in the detector forward region. In contrast, at electron-positron colliders with center-of-mass energies ($\sqrt{s}$) around 10 GeV or lower, such as the BelleII [2, 3] and BESIII [4, 5] experiments, the particle multiplicity is typically below 10, and their detectors usually emphasize on the PID performance, achieving 1-1 correspondence could be much easier. Our study focuses on the electron-positron Higgs factory [6, 7, 8, 9, 10] which is regarded as the highest-priority future collider as it offers excellent opportunity to discover new physics [11, 12], especially via precision measurements of the Higgs boson. It operates at center-of-mass energies ranging from 91.2 GeV to several TeV, typically producing $\mathcal{O}(100)$ visible particles distributed in narrow jets in hadronic events.
The concept of 1-1 correspondence evolves from the Particle Flow Algorithm (PFA) [13, 14, 15, 16], with the key idea to trace each individual particle, classify and associate corresponding detector hits, and measure the particle 4-momentum using the most suitable sub-detector system. Originating from the ALEPH [16] experiment, the PFA has become the guiding principle for multiple conceptual detector designs at electron-positron Higgs factories [9, 10, 17] and detector upgrade projects at HL-LHC [18]. Benefit from the rapid development of artificial intelligence, many machine learning algorithms have been implemented to enhance the PFA performance [19, 20, 21]. The current PFA performance is primarily limited by the confusion [22], which includes the effects of fake particles, failures in track-cluster matching, particle loss due to shower overlap, etc. These confusion effects violate the 1-1 correspondence relation. Therefore, the goal of 1-1 correspondence reconstruction is to approach confusion-free particle flow reconstruction and to identify the species of all visible particles.
Refer to caption Refer to caption
Figure 1: AURORA detector geometry and an event display of $\nu\bar{\nu}H,H\to gg$ ($\sqrt{s}$ = 240 GeV) event.
The performance of PFA can be quantified using the Boson Mass Resolution (BMR) [9, 23], which represents the relative mass resolution of massive bosons (such as Higgs, $Z$, and $W$ bosons) decaying into hadronic final states. At the electron-positron Higgs factory, a BMR smaller than 4% is required to separate the $q\bar{q}H$ signal from $q\bar{q}Z$ background using the recoil mass method [9, 24]. A smaller BMR is beneficial for all physics measurements with hadronic final states, especially for flavor physics measurements and new physics searches since they rely on accurate measurements of missing energy and momentum. Taking reference to the Circular Electron-Positron Collider (CEPC) Conceptual Design Report (CDR) [9], its baseline detector design achieves a BMR of 3.7% [25, 23], meeting the requirements for Higgs measurements. In terms of PID, the CEPC baseline detector demonstrates a typical efficiency of 99.5% and a misidentification rate around 1% for isolated lepton identification [26, 27]. It can also distinguish charged hadrons ($\pi^{\pm}$, $K^{\pm}$, $p/\bar{p}$) using techniques like ${\rm d}E/{\rm d}x$ and time-of-flight (TOF) measurements [28, 29].
To pursue 1-1 correspondence reconstruction, we develop the AURORA (ApparatUs for RecOnstRuction with Advanced algorithm) detector concept and PROOF (Particle Reconstruction with One-to-One correspondence at Higgs Factory) reconstruction framework, based on the CEPC CDR baseline design [9]. AURORA features a high-granularity 5-dimensional (5D) calorimetry as its core, with a unified time resolution of 100 ps for each calorimeter hit. The left panel of Fig. 1 shows a cut-away view of AURORA, with more details provided in Methods. Combining the CEPC baseline PFA Arbor [14] and Particle Transformer (ParT) [30], PROOF performs a universal identification of all reconstructed particles, simultaneously distinguishing different confusion types and identifying particle species. ParT is a particle cloud- [31] and Transformer-based [32] machine learning model adapted for high-energy physics experiments, which has been applied in jet flavor tagging at the LHC [33, 34, 35] and jet origin identification (JOI) at the electron-positron Higgs factory [36]. Following the convention of CEPC performance studies [9, 25, 37], we simulate 1 million $\nu\bar{\nu}H,H\to gg$ ($\sqrt{s}$ = 240 GeV) events with AURORA detector and reconstruct these events with Arbor. A reconstructed $\nu\bar{\nu}H,H\to gg$ event is displayed in the right panel of Fig. 1, and more details on simulation and reconstruction are described in Methods.
Refer to caption Refer to caption
Figure 2: Left: Schematic diagram of visible-reconstructed particle mapping. “Charged without cluster/track” means a charged particle does not create enough hits to form a cluster/track (mainly due to limited detector acceptance and efficiency), while “Charged with cluster lost” refers to cases where both a cluster and track are formed but do not match each other, resulting in fake particles. Right: Fraction of truth visible energy mapped into different categories.
Using truth-level information in the Monte Carlo (MC) simulation, we establish a mapping between visible particles and reconstructed ones, as illustrated in the left panel of Fig. 2. The mapping consists of three main categories:
  • 1-1 correspondence (solid arrows in Fig. 2): one visible particle is mapped to only one reconstructed particle. This category consists of three types, represented in shades of green in Fig. 2. One type is well-reconstructed, where a cluster is a must for neutral particles, and both track and cluster are required for charged ones. The other two types contain charged particles that generate only track or cluster, mainly caused by the limitations of detector acceptance (e.g. in the forward region) and efficiency (e.g. dead zone or energy/momentum measurement thresholds).
  • Fake particles (in orange and yellow in Fig. 2): occur when one visible particle is mapped to multiple reconstructed particles. Due to detector inefficiency and particle interactions with upstream materials, the calorimeter hits generated by a single particle may be grouped into multiple clusters. For a charged particle, its track could be matched to only part of or even none of these clusters, while the remaining clusters are reconstructed as nearby extra neutral particles. These neutral particles that originate from charged particle shower fragments and un-associated clusters, represented in orange in Fig. 2, can cause a double-counting in the total reconstructed energy. For simplicity, these neutral particles are called fake particles in this study.
  • Lost particles (in gray in Fig. 2): visible particles with no corresponding reconstructed particles. Particles can be lost due to limited detector acceptance [^1] In this article, we define visible particles as those that create detector hits. On the other hand, the limited detector acceptance causes particle loss, i.e., AURORA has an overall geometry acceptance of $|\cos\theta|\approx 0.99$, where it cannot record particles traveling in the forward direction. Since particle loss due to detector acceptance certainly impacts the overall physics measurement, we list these particles in dashed boxes in Fig. 2 to complete the picture and also consider them when calculating visible energy fractions in the following text., selection criteria, and shower merging. Selection criteria are applied to improve the reconstruction performance, such as vetoing fake particles, but inevitably they could exclude genuine particles, usually low-energy ones. Shower merging (dotted arrow in Fig. 2) is another major cause of particle loss, particularly within high-energy jets where nearby or overlapping calorimeter showers can be merged into one cluster, especially at limited calorimeter granularity. For instance, the reconstruction of $\pi^{0}\to\gamma\gamma$ can be particularly sensitive to shower merging.
According to this mapping with truth information, we label reconstructed particles into 15 types, including 10 types of visible particles ($e^{\pm}$, $\mu^{\pm}$, $\pi^{\pm}$, $K^{\pm}$, $p/\bar{p}$, $\gamma$, $K_{L}^{0}$, $n$, $\bar{n}$, and others including $K_{S}^{0}$ and $\Lambda$ that not yet decay within the tracker volume) and 5 types of confusion (such as different kinds of fake particles, charged particle without track, etc). Each reconstructed particle is characterized by 55 observable variables as inputs to ParT. More details can be found in Methods.
The quantitative analysis [38] shows that fake particles are the leading contribution to the BMR. In our simulation, on average, a $\nu\bar{\nu}H,H\to gg$ event has 80 visible particles with a total energy of 135 GeV at the truth level, while the reconstructed fake particles lead to a double-counted energy of 6 GeV, severely degrading the BMR. PROOF can identify the fake particles with a typical efficiency of 77% and a purity of 97.5%, consequently suppressing the average double-counted energy from fake particles to 0.7 GeV (by nearly one order of magnitude) at the minor cost of mis-excluding visible particles with a total energy on the order of 0.5 GeV. Consequently, PROOF achieves a BMR of 2.75% (see the left panel of Fig. 3), where the effects of fake and mis-excluded particles get balanced. Compared to the CEPC CDR [39], the BMR is improved by 25%, where 10% is from the detector geometry optimization to AURORA [40], and 15% is from the fake particle suppression. The contributions from the three categories of visible-reconstructed particle mapping to the BMR are not proportional to their corresponding energy fractions. The contribution from the 1-1 correspondence category primarily stems from the detector resolution, which is orders of magnitude lower compared to the total energy of particles within this category. Meanwhile, the contributions from the fake and lost categories are comparable to their corresponding total energies. We found that the 1-1 correspondence category contributes to half of the BMR, while the fake and lost categories each contribute a quarter, combined quadratically.
Refer to caption Refer to caption
Figure 3: Left: Invariant mass distributions of hadronically decayed Higgs, $W$, and $Z$ bosons derived by the combination of AURORA and PROOF to approach 1-1 correspondence reconstruction. Right: Confusion matrix of well-reconstructed particles identification. Both are after removing identified fake particles.
Referring to the truth level energy of visible particles, we calculate the fraction of total visible energy corresponding to different categories of the mapping. Ignoring the remaining fake particles [^2] The remaining fake particles could violate the 1-1 correspondence relationship, as one visible particle could be mapped to one leading reconstructed particle together with one or even several fake particles. However, as the energy of the fake particle is usually much lower than the leading one, the fake particle only has a tiny impact on the reconstruction and identification of the original visible particle. Meanwhile, the total energy of remaining fake particles only contributes to a sub-percentage level of total visible energy. Therefore, we ignore these remaining fake particles in our study., we observe that over 90% of the total visible energy is mapped to the well-reconstructed particles, 4% to charged particles without clusters (with over half of this due to calorimeter acceptance/inefficiency), 2.5% to charged particles without tracks (mostly in the forward region), and 2.6% of the energy is lost due to limited detector acceptance, shower merging, and mis-excluded genuine particles, as illustrated in the pie chart in Fig. 2. The detector acceptance dominates the confusion of charged particles without clusters or tracks. Summing the energy fraction of well-reconstructed particles and confusion type purely originating from detector acceptance, we consider that nearly 95% of the total visible energy maintains the 1-1 correspondence relationship.
Focusing on the well-reconstructed particles of 1-1 correspondence type, the PID performance of PROOF is shown in the confusion matrix in the right panel of Fig. 3. The identification efficiencies (diagonal matrix elements) range from 97% to nearly 100% for charged particles and photons, and from 75% to 80% for three types of neutral hadrons. Since the PID performance depends on particle kinematics, we extract the traces (i.e. sums of diagonal elements) of sub-matrices in different particle momentum and polar angle ranges, as shown in Fig. 4. For photons and charged particles, a high-trace plateau is observed within the momentum range of 0.8 to 20 GeV and $|\cos\theta|<0.96$. Inefficiencies primarily arise due to limited cluster-level information in the low-momentum region and degraded PID performance in high-momentum and detector-forward regions. Similar behavior is observed for neutral hadrons, which shows that the trace is approaching 3 in a limited phase space, indicating that even neutral hadrons can be efficiently separated. These observations lead to the conclusion that 1-1 correspondence is feasible at future electron-positron Higgs factories.
It should be remarked that many simplifications are made to facilitate the analysis. To truly realize 1-1 correspondence reconstruction, dedicated studies are required in the future. For instance, quantifying sub-detector performance specifications—such as the time (TOF) resolution of calorimeters, ${\rm d}E/{\rm d}x$ resolution of gaseous trackers, and detector geometry optimization—is essential. The reconstruction performance of the primary vertex is also relevant for TOF measurements. The impact of beam-induced background needs to be quantified. Modeling and verifying detailed detector responses are crucial for controlling systematic uncertainties of MC simulations. Secondary particles, treated in the same manner as primary particles in this study, can significantly affect reconstruction performance and physics measurements. For example, electron-positron pairs from photon conversion and muons from hadron decays could contribute to the background in measurements that depend on (semi-)leptonic decays. Therefore, these secondary particles require dedicated identification methods, which could be developed within the future PROOF framework.
Figure 4: Traces of $6\times 6$ ($e^{\pm}$, $\mu^{\pm}$, $\pi^{\pm}$, $K^{\pm}$, $p/\bar{p}$, $\gamma$) (left) and $3\times 3$ ($K_{L}^{0}$, $n$, $\bar{n}$) (right) sub-matrices as function of momentum and polar angle $|\cos\theta|$.
Prospects
Our study demonstrates the feasibility of 1-1 correspondence reconstruction at the electron-positron Higgs factory, it can significantly improve the BMR performance, and simultaneously identify 10 different kinds of visible particles. Meanwhile, the performance bottleneck of 1-1 correspondence in our current setup is identified to be the detector acceptance and neutral hadron identification, which could be ameliorated via dedicated algorithms and detector optimization. This section discusses the prospects for high-level reconstruction and physics measurements with an ideal 1-1 correspondence reconstruction.
Refer to caption Refer to caption
Figure 5: Top left: Neutral hadron energy resolution using hadronic calorimeter (HCAL, blue and green lines) or TOF (red and yellow lines). Top right: Jet origin identification performance (characterized by flavor tagging efficiencies and charge flip rates) in two scenarios: Scenario 1 (S1) assumes perfect identification of leptons and charged hadrons, while Scenario 2 (S2) extends S1 by further assuming perfect identification of $K^{0}_{L}/K^{0}_{S}$, $n/\bar{n}$, and Λ/$\Lambda/\bar{\Lambda}$. Bottom: Anticipated upper limits on branching ratios of Higgs exotic and rare decays with different performances of JOI and BMR.
Since 1-1 correspondence reconstruction provides the particle mass information via PID, the energy of visible particles can be determined by the TOF. 5D calorimetry provides TOF measurements for both charged and neutral particles. For charged particles, their energies/momenta can be determined by three methods: shower energy via calorimeter, track momentum via tracker, and TOF via calorimeter, while the TOF method has a comparative advantage in the forward region. For neutral particles, their energies can be determined by both the shower energy and the TOF via calorimeter, with TOF showing a significant advantage for low-energy neutral hadrons, as demonstrated in the top left panel of Fig. 5. Therefore, by combining these measurements of particle energy, such as selecting the most suitable method according to the particle species and their kinematics, we can significantly improve the energy/momentum resolution of different particles, and consequently improve the BMR. Considering the improvement in particle energy measurement can also benefit pattern recognition and confusion control, we estimate that this approach could further enhance the BMR by 15–20%, meaning that a BMR of 2.2–2.4% could, in principle, be achieved.
1-1 correspondence reconstruction can also significantly enhance the JOI performance. JOI is the procedure to determine the type of quark/gluon (11 types are typically considered: $b$, $\bar{b}$, $c$, $\bar{c}$, $s$, $\bar{s}$, $u$, $\bar{u}$, $d$, $\bar{d}$, $g$) from which a jet originates. In a recent study [36], we realize JOI with the CEPC CDR baseline detector. It can simultaneously identify $b$, $c$, $s$ quarks with efficiencies of 70–90% and $u$, $d$ quarks with efficiencies of 40%, meanwhile, the charge flip rates of quarks and anti-quarks are controlled to be 10–20%, with an ideal lepton and charged hadron identification. JOI significantly boosts the accuracy of flavor-sensitive physics measurements, such as $H\to s\bar{s}$ and $H\to sb$ decays [36], $|V_{cb}|$ [41], $\phi_{s}$ [42], and weak mixing angle [43] measurements. 1-1 correspondence provides the identification of $K^{0}_{L}$, $n$, and $\bar{n}$, while in principle the $K^{0}_{S}$ and $\Lambda$ could also be identified [44], therefore, providing the PID information of all these neutral hadrons, we observe a significant improvement on JOI performance especially the identification performance of $u/d/s$ quarks.
BMR and JOI are key performances of event reconstruction at electron-positron Higgs factories. Their improvements will benefit most physics measurements with hadronic final states. For example, the Higgs to invisible final states is a key portal to detect dark matter at Higgs factories. The original upper limit at 95% confidence level is quantified to be $\mathcal{O}(0.1\%)$ [45] with a BMR of 3.7%. Improving the BMR to 2.75%, or even 2.2%, will consequently enhance the upper limit by 10–12% [46]. Higgs exotic and rare hadronic decays are also sensitive probes to new physics. Using JOI technology, these decay modes can be typically limited to $10^{-4}$$10^{-3}$ using $\nu\bar{\nu}H$ and $\ell^{-}\ell^{+}H$ processes [36]. With the improvement in JOI alone, the constraints could be enhanced by 5–40%. Considering also the improved BMR, the sensitivities are expected to be further enhanced by another 5–10%, see the bottom panel of Fig. 5. In short, 1-1 correspondence could significantly boost the discovery power of electron-positron Higgs factories.
Based on this study, we are confident that 1-1 correspondence reconstruction will be a new paradigm for event reconstruction at the high-energy frontier. For the electron-positron Higgs factory, 1-1 correspondence reconstruction is a goal that should be pursued, as it can significantly enhance the physics reach and discovery power. It is also a goal that could be achieved—presumably very soon—through integrating innovative detector technologies, advanced reconstruction algorithms, and artificial intelligence.
Acknowledgements
We would like to express our sincere gratitude to Jianchun Wang for his continuous support throughout this work. We also thank Hengyu Wang for his study on HCAL resolution, which provided an important reference for our research. This work is supported in part by the National Natural Science Foundation of China under grant No. 12042507, the National Key R&D Program of China under Contracts No. 2022YFE0116900, and the National Key Program for S&T Research and Development (2023YFA1606300). Additional support is provided by the Fundamental Research Funds for the Central Universities, Peking University.